The roof of a lean-to-garden shed has a gradient of 0.35.find the height....they have given the base as 1.5m and a shorter height 1.8m its like a house
Answer by stanbon(75887)

The roof of a lean-to-garden shed has a gradient of 0.35.find the height....they have given the base as 1.5m and a shorter height 1.8m its like a house
-------
It sounds like you have a right triangle with base = 1.5,
----
Use a proportion to get:
(height of that triangle)/1.5 = 35/100
height of that triangle = 1.5*0.35 = 0.5250
-----
Ans: greater height of the lean-to = 1.8 + 0.5250 = 2.325 meters
-----------------
Cheers,
Stan H.